HIGH RESOLUTION SPECTROSCOPY AND AB INITIO CALCULATIONS ON HfCl

R. S. RAM, Department of Chemistry, University of Arizona, Tucson, AZ 85721; A. G. ADAM, Department of Chemistry University of New Brunswick, Fredericton, N.B., Canada E3B 6E2; A. TSOULI, J. LIÉVIN, Université Libre de Bruxelles, Laboratoire de Chimie Physique Moléculaire, CP 160/09, Av. F. D. Roosevelt 50, Bruxelles, Belgium; \underline{P. F. BERNATH}, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1.